OpenStudy (anonymous):

Find the base edge of a square prism with a hight of 5m and a lateral area of 100 meters squared.

OpenStudy (anonymous):

Well the area of the square will be 25 so we will have left 75 sq. meters the r of the circle will be \[\sqrt{5}\] so each side will have length \[\pi \sqrt{5}\div\]2